OpenStudy (anonymous):

I am getting confused with this chapter about notes and bank discounts, does anyone know about this? Thanks The Treasury Department auctioned $21 billion in 3-month bills in denominations of $10,000 at a discount rate of 4.965%. What would be the effective rate of interest? Round your answer to the nearest hundredth percent.

OpenStudy (phi):

I think you pay less than face value for the 3-month bill. To figure out the amount you pay, first change the yearly discount rate to quarterly (divide by 4... because 3 months is ¼ of a year) 0.04965/4 (1- 0.04965/4)=0.9875875 the amount you pay for $1 multiply by the face value of the bill 10,000*0.9875875= 9,875.875 to find the effective interest rate, find the amount of interest= 10000-9,875.875 and divide the interest by 9,875.875 (10000-9,875.875)/9,875.875= 0.0125685065880238 multiply by 4 to get the effective yearly rate 0.0125685065880238*4= 0.0502740263520952 in percent, that is 5.0274 % rounded to the nearest hundredth we get 5.03%